OpenStudy (anonymous):

10m-9n=15, 5m-4n-10 I need to find a solution or elimination

OpenStudy (cwrw238):

10m-9n=15, 5m-4n = 10 multiply 2nd by 2 10m - 8n = 10 oh - should it b, 5m-4n=10 not 5m-4n -10?

OpenStudy (anonymous):

it's = not -

OpenStudy (cwrw238):

ok so if we subtract the last equation from the first we eliminate m -n = -5 n = 5 plug n=5 into first equation 10m - 9(5) =15 10m = 14+45 = 60 m = 6 solution is m=6,n=5

OpenStudy (cwrw238):

oh - theres a typo on the 4th line in my first post correct is 10m - 8n = 20 this doesn't affect the solution

OpenStudy (cwrw238):

you ok with this?

OpenStudy (anonymous):

yes sorry was typing a paper, and thank you for the help :)

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

Can you help me graph y=2/3x+1 and y=-1+8/12x in the same graph